As the customer service analyst, you have documentation for 126 service calls concerning incorrect product delivery from your study of 30 samples. What is the upper control limit for this control chart? (Keep two decimal places in your answer)
Answer #1

Answer: UCLc= 10.35

Explanation: We will use C chart

Incorrect product delivery= total defects= 126

Total samples= 30

Mean of defects, C=  total defects/total samples= 126/30= 4.2

C-Chart using z= 3
Mean, C= AVERAGE(all defects) 4.20
Standard Dev (STD)= sqrt (mean) 2.05
UCL= C + z*STD 10.35

Please ask, if you have any doubts through the comment section. Do rate the answer